# LeTO: Learning Constrained Visuomotor Policy with Differentiable Trajectory Optimization

### 30 Jan 2024 v1

@@ V-A Move-the-stack
For imitation learning algorithms, the task is challenging because: 1) The policy needs to generate very smooth trajectories. Any roughness in the trajectory can lead to vibrations at the robot’s end-effector, causing the cups to shake. This vibration can accumulate and lead to the chopstick dropping. 2) Since a chopstick is introduced during policy rollout, this out-of-distribution (OOD) observation adds greater uncertainty to the network’s input. This uncertainty can make the network’s output jerkier, resulting in a poorer quality trajectory.
The results are shown in Table IV. From the results, LeTO excels in terms of trajectory smoothness and, correspondingly, also boasts the highest success rate. This demonstrates that due to the presence of end-to-end trajectory optimization, LeTO can robustly maintain its effectiveness even when facing OOD data. Another point worth noting is that when chopsticks fall (which is also a type of OOD data), both the diffusion policy and LeTO can still move the stacked cups and place them on the black board. In contrast, LSTM-GMM often exhibits erroneous behaviors, such as moving to incorrect locations or acting erratically (see the attached video). This showcases the robustness of LeTO and the diffusion policy.
V-B Arrange-chopsticks
The challenges of this task include: 1. It is a long-horizon task, requiring the policy to have the capability to represent complex tasks. 2. Due to the slender nature of chopsticks, there is a high risk of the robot making forceful contact with the table when attempting to grasp them. For the sake of hardware safety, such collisions should be minimized as much as possible.
The results are shown in Table V. Both LeTO and diffusion policy have high success rate, but the contact forces in LeTO tests are significantly smaller.
VI Discussion and future work
VI-A Hyperparameter, Training Time, and Training Stability
@@
Table VI compares the training times of LSTM-GMM, diffusion policy, and LeTO. We acknowledge that one limitation of LeTO is the computational overhead introduced by the differentiable optimization layer. Future work could look into improving the computational efficiency of the differentiable optimization process, potentially through the use of more efficient solvers or by learning approximations of the optimization layer.
@@ VI-B LeTO in Low-data and High-data Regimes
Furthermore, exploring policies that excel in low-data regimes based on differentiable optimization layers is an intriguing research direction. This direction may not be limited to imitation learning; for instance, reinforcement learning could also benefit from similar concepts. By constructing differentiable optimization layers using prior knowledge of the model, policy may reduce reliance on data.
VI-C Multi-modal Distribution Modeling of Differentiable Optimization
In this section, we will demonstrate LeTO is an implicit model. The forward pass of the DTO layer during rollout is as follows
$\hat{\mathbf{y}}^{\star}_{t}=\underset{\hat{\mathbf{y}}_{t}}{\operatorname{argmin}}$ $\frac{1}{2}\hat{\mathbf{y}}_{t}^{\mathrm{T}}\bar{Q}\hat{\mathbf{y}}_{t}+e_{t}^{\mathrm{T}}\hat{\mathbf{y}}_{t},$ (24)
[…]
$a_{\text{min}}\Delta t\leq\hat{v}_{t}-\hat{v}_{t-1}\leq a_{\text{max}}\Delta t.$ (25)
In equation (24), $\bar{Q}$ is constant after training and $e_{t}$ is a function with the observation and hidden state which is related to past observations. Note that the observation only consists of the robot pose and the camera images. However, the trajectory optimization not only works on robot pose, but also on velocity and acceleration. As shown in equation (25), $\hat{v}_{t-1}$ is the last executed velocity command and it is not included in the observation and the optimization variable $\hat{\mathbf{y}}_{t}$. Even under similar observations (which means $e_{t}$ does not change too much and so does the cost function), $\hat{v}_{t-1}$ can be very different. Since the outputs of LeTO are predicted action trajectories, variations in initial constraints $a_{\text{min}}\Delta t+\hat{v}_{t-1}\leq\hat{v}_{t}\leq a_{\text{max}}\Delta t+\hat{v}_{t-1}$ can lead to entirely different predicted trajectories. Thus, in this way, LeTO is an implicit model and can generate very different action outputs under similar observations.
While we demonstrate that LeTO is implicit, we have yet to provide a rigorous proof that LeTO can model multi-modal distributions, or demonstrate the extent to which it is capable of modeling multi-modal distributions. We believe that mathematically proving that LeTO, along with similar implicit differentiable optimization models, can capture multi-modal distributions, and developing imitation learning algorithms that effective model multi-modal distribution while maintaining interpretability and safety by integrating differentiable optimization and generative models, represents highly promising research directions.
VII acknowledgements
This work was partially supported by the United States Department of Agriculture (USDA; grant number 2023-67021-39072). This article solely reflects the opinions and conclusions of its authors and not USDA.

### 18 Mar 2024 v2

@@ VI-A Hyperparameter, Training Time, and Training Stability
Table VIII compares the training times of LSTM-GMM, diffusion policy, and LeTO. We acknowledge that one limitation of LeTO is the computational overhead introduced by the differentiable optimization layer. Future work could look into improving the computational efficiency of the differentiable optimization process, potentially through the use of more efficient solvers or by learning approximations of the optimization layer.
@@ VI-B LeTO in Low-data and High-data Regimes
Furthermore, exploring policies that excel in low-data regimes based on differentiable optimization layers is an intriguing research direction. By constructing differentiable optimization layers using prior knowledge of the model, policy may reduce reliance on data.
VI-C LeTO in Reinforcement Learning
Due to the differentiability, the DTO layer can be designed to integrated into the robot learning pipelines, enabling end-to-end optimization and inference. This opens up the possibility of extending LeTO to reinforcement learning, offering a promising future research direction. By integrating differentiable trajectory optimization with reinforcement learning, it is possible to ensure that the generated actions comply with the designed constraints, thereby significantly enhancing the safety and interpretability of the policy.
VI-D LeTO in High-safety Scenarios
The safe and constraint-controlled output form of LeTO holds significant potential in robot learning scenarios where safety is of utmost importance, such as in human-robot interaction and surgical robotics. Investigating the application of LeTO, or more broadly, robot policies that integrate differential optimization in these settings, represents a highly interesting and promising research direction.